Questions 1 through 6 are based on the information provided below. The sales, income from operations, and invested assets for Johnson Company are as follows:







Chapter 24 Homework Question 1: (0.5 point) Required: Using the Dupont formula, calculate the profit margin for Division E. When entering the answers in Blackboard, omit $ signs. Round all calculations to one decimal place.

Chapter 24 Homework Question 2: (0.5 point) Required: Using the Dupont formula, calculate the investment turnover for Division E. When entering the answers in Blackboard, omit $ signs. Round all calculations to one decimal place.

Chapter 24 Homework Question 3: (0.5 point) Required: Using the Dupont formula, calculate the rate of return on investment for Division E. When entering the answers in Blackboard, omit $ signs. Round all calculations to one decimal place.

Chapter 24 Homework Question 4: (0.5 point) Required: Using the Dupont formula, calculate the profit margin for Division F. When entering the answers in Blackboard, omit $ signs. Round all calculations to one decimal place.

Chapter 24 Homework Question 5: (0.5 point) Required: Using the Dupont formula, calculate the investment turnover for Division F. When entering the answers in Blackboard, omit $ signs. Round all calculations to one decimal place.

Chapter 24 Homework Question 6: (0.5 point) Required: Using the Dupont formula, calculate the rate of return on investment for Division F. When entering the answers in Blackboard, omit $ signs. Round all calculations to one decimal place.










Sales Income from operations
Invested Assets
Division E $4,000,000 $552,000 $2,500,000 Division F 4,800,000 600,000 1,920,000
Answer #1

Question 1: (0.5 point) Required: Using the Dupont formula, calculate the profit margin for Division E.

Profit margin= Income from operations/Sales

Division E= $552000*100/4000000= 13.8%

Question 2: (0.5 point) Required: Using the Dupont formula, calculate the investment turnover for Division E.

Investment turnover= Sales/Invested assets

= $4000000/2500000= 1.6

Question 3: (0.5 point) Required: Using the Dupont formula, calculate the rate of return on investment for Division E.

Rate of return on investment= Profit margin*Investment turnover

= 13.8%*1.6= 22.1%

Question 4: (0.5 point) Required: Using the Dupont formula, calculate the profit margin for Division F.

Profit margin= Income from operations/Sales

Division F= $600000*100/4800000= 12.5%

Question 5: (0.5 point) Required: Using the Dupont formula, calculate the investment turnover for Division F.

Investment turnover= Sales/Invested assets

= $4800000/1920000= 2.5

Question 6: (0.5 point) Required: Using the Dupont formula, calculate the rate of return on investment for Division F.

Rate of return on investment= Profit margin*Investment turnover

= 12.5%*2.5= 31.3%